OpenStudy (anonymous):

how can i do the derivates of (x+1)^3/x3/2

OpenStudy (anonymous):

Is this the function? \[(x + 1)^3/x ^{3/2}\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Ok... rewrite: \[y = (x + 1)^3x ^{-3/2}\] Product rule: \[y'= 3(x + 1)^2(1)x ^{-3/2}+(x+1)^3(-3/2)x ^{-5/2}\] \[y'=(x+1)^2x ^{-3/2}[3+(-3/2)(x+1)x ^{-1}]\] \[y'=(x+1)^2[3/2-3/(2x)]/x ^{3/2}\] \[y'=(x+1)^2[(3x-3)/(2x)]/x ^{3/2}\] \[y'=3(x-1)(x+1)^2/(2x ^{5/2})\] Someone else may wanna check this over, though...

OpenStudy (anonymous):

Yup I checked. It's definitely right.

OpenStudy (anonymous):

i got lost on the second line

OpenStudy (anonymous):

On the second line I factored out (x + 1)^2 and x^(-3/2).